# Neural variational inference for cutting feedback during uncertainty propagation

### v1

@@ Abstract
In many scientific applications, uncertainty of estimates from an earlier (upstream) analysis needs to be propagated in subsequent (downstream) analysis, without feedback. Cutting feedback methods, also termed cut-Bayes, achieve this by constructing a cut-posterior distribution that prevents backward information flow, facilitating unidirectional uncertainty propagation. However, cutting feedback using sampling methods like nested MCMC is computationally challenging. Recently variational inference (VI) methods have been developed to approximate the cut-posterior. However, they need two variational approximations and require access to the upstream data and model. In this manuscript we propose, NeVI-Cut, a provably accurate and modular neural network-based variational inference method for cutting feedback. We directly utilize samples from the upstream analysis without requiring access to the upstream data or model. This simultaneously preserves modularity of analysis and reduces approximation errors by avoiding a variational approximation for the upstream model. Conditioning on each upstream sample, we then use normalizing flows (neural network-based generative models) to specify the conditional variational family for the downstream parameters. We estimate the conditional cut-posterior by optimizing the Monte Carlo average of the variational loss over all the upstream samples. We provide theoretical guarantees on the NeVI-Cut estimate to approximate any cut-posterior. Our results are in a fixed-data regime. We do not assume any asymptotic paradigm or even sample size growth, and go beyond just establishing universal expressiveness of normalizing flows. We provide convergence rates of the actual variational solution and quantify how richness of the neural architecture and the complexity of the target cut-posterior dictate the approximation quality. In the process, we establish new results on uniform Kullback-Leibler approximation rates of conditional distributions by common flow classes like Unconstrained Monotonic Neural Network and neural Rational Quadratic Splines. A triply stochastic scalable algorithm is devised to implement the method. Simulation studies and two real-world analyses illustrate how NeVI-Cut achieves significant computational gains over traditional cutting feedback methods and is considerably more accurate than parametric variational cut approaches.
Keywords: Bayesian analysis, Cutting Feedback; Modular Inference; Variational Inference; Normalizing Flows; Neural Networks, Universal approximation rates.
@@ 1 Introduction
Across many scientific fields, estimates of quantities from an upstream analysis are prestored to be reused in many downstream analyses. Often these upstream parameter estimates are quantified with uncertainty (Monte Carlo samples, ensemble members, or Bayesian posterior samples) and any use of these parameters in any downstream analysis mandates proper propagation of uncertainty. For example, downstream climate impacts studies (e.g., Butler et al.,, 2022; Park et al.,, 2023; Rao et al.,, 2024) commonly use ensemble outputs from climate models such as CMIP6 (Thrasher et al.,, 2022) as probabilistic projection samples, rather than using single climate trajectories, so that the spread from the climate model propagates into sectoral outcomes. Similarly, in environmental health and air quality research, widely used data products explicitly quantify uncertainty. The Global Burden of Disease project releases fine particulate matter (PM2.5) exposure and risk estimates as $1000$ posterior draws (Institute for Health Metrics and Evaluation,, 2024) that are used downstream in attributable-burden calculations (Murray et al.,, 2020). Other widely used high-resolution exposure surface data products also provide estimates with uncertainty (PM2.5 and NO2 estimates developed in Di et al., 2019b, ; Di et al., 2019a, ). Importance of incorporating this uncertainty in downstream health outcomes analysis has been routinely emphasized in the related literature on exposure measurement error (Peng and Bell,, 2010; Chang et al.,, 2011; Spiegelman,, 2016; Comess et al.,, 2024). Other examples of uncertainty quantified prestored outputs include posterior distributions of cosmological parameters, termed as Planck MCMC chains, posterior samples of confusion matrices of cause-of-death classifiers (Pramanik et al., 2025a, ), and probabilistic word-embeddings used in large language models (LLM) like word2Gauss and Bayesian skip-gram (Vilnis and McCallum,, 2014; Bražinskas et al.,, 2018).
@@
In many of the aforementioned applications, when propagating uncertainty to a downstream Bayesian analysis, it is desirable not to have feedback, which is the statistical phenomenon of the downstream data informing an updated distribution for the upstream parameters. This is because the upstream quantities are often more fundamental and causally precedent (e.g., air pollution concentrations or climate data) than downstream ones (e.g., individual-level health outcomes). Additionally, downstream models are often misspecified for the upstream quantities. For example, in downstream health association studies, upstream estimates of area-level air pollution outputs are often used only as proxies for the true unobserved individual-level exposures (Zeger et al.,, 2000). Similarly, confusion matrix estimates of classifiers trained on specific populations are often used in downstream calibration of different target populations (Pramanik et al., 2025a, ). Hence, it is undesirable to update these upstream parameter distributions based on the downstream data and model that are subject to model misspecifications, dataset shifts, and biases. Furthermore, many of these upstream quantities (e.g., climate outputs, air pollution estimates) are used in many different downstream analyses, and it is impractical to expect revised upstream posteriors to be published after every downstream study conducted by different groups. In such settings, when conducting a downstream Bayesian analysis using uncertainty-quantified upstream quantities, one should prevent feedback and preserve the original upstream posterior while coherently propagating its uncertainty to the downstream quantities.
Feedback occurs in sequential Bayesian analysis if the upstream parameter distribution is used as a prior for that parameter in the downstream analysis (e.g., the analysis in Pramanik et al., 2025b, ) or if the upstream samples are used downstream using recursive Bayes methods (see e.g., Taylor et al.,, 2025, for a recent work). Both of these approaches approximate a full Bayesian analysis using both upstream and downstream data, which allows information sharing in all directions to infer all parameters jointly from all data, thereby resulting in feedback. The idea of explicitly stopping downstream to upstream feedback is known as Cutting feedback (Liu et al.,, 2009; Plummer,, 2015; Jacob et al.,, 2017). Instead of trying to estimate the conventional Bayesian posterior, cutting feedback methods target the cut-posterior, a modification that yields a valid joint distribution over all parameters while preserving the distribution of the upstream parameters from the first analysis.
There is now a large and rapidly growing literature on cutting feedback and related methods because of their broad applicability. Cutting feedback methods has been applied in various areas, including pharmacology (Lunn et al.,, 2009), economics (Murphy and Topel,, 2002), environmental health (Peng and Bell,, 2010, their two-stage Bayesian model is a version of nested MCMC used in cut models, although they didn’t make the explicit connection), and causal inference (Zigler,, 2016). Moss and Rousseau, (2024) use cut-posteriors in a hidden Markov model, preventing feedback from the emission distribution to transition parameter estimation. Smith et al., (2025) propose cutting feedback methods for misspecified copula models, improving accuracy compared to conventional Bayesian posteriors. Carmona and Nicholls, (2020, 2022) introduced semi-modular inference, that interpolates posteriors between Bayes and cut-Bayes, allowing partial feedback regulated by a hyperparameter for greater flexibility. More generally, cut-Bayes can be viewed as a case of modular Bayesian inference involving several data-model pairs where certain directions of information flow are unwarranted. Liu and Goudie, (2025) present a general framework for this, formally defining modules via Markovian relationships or directed acyclic graphs, and developing methods for general modular inference. We refer to Nott et al., (2023) for an overview of recent developments in Bayesian inference for misspecified models, including cutting feedback methods.
@@
A major thread of developments on cutting feedback methods has been computational. The cut function in the OpenBUGS package modifies the MCMC conditional updates to restrict feedback between modules; however, Plummer, (2015) showed that this algorithm does not converge to the target cut-posterior. In general, designing a Markov Chain Monte Carlo (MCMC) that has the cut-posterior as the invariant distribution can be challenging. A principled alternative is the multiple imputation-based nested MCMC. This involves first running an MCMC for the upstream analysis to obtain a set of posterior samples, and then for each of these samples, running an MCMC for the downstream analysis, and then combining over all the MCMC posteriors. The repeated MCMC runs for the downstream analysis significantly increase computational cost. Plummer, (2015) proposed a tempered cut algorithm for approximating the cut-posterior. Jacob et al., (2020) developed an unbiased MCMC estimator using coupled chains and demonstrated the use for approximating the cut-posterior. Liu and Goudie, (2022) introduced a stochastic approximation adaptive MCMC method to approximate the cut-posterior, improving convergence and stability compared with the openBUGS algorithm and tempered cut. Pompe and Jacob, (2021) designed a posterior bootstrap method for modular inference, providing valid uncertainty quantification. Chakraborty et al., (2023) introduce a likelihood-free approach using Gaussian mixtures to approximate cut and semi-modular posteriors when the likelihood is intractable. Hutchings et al., (2025) proposed the Emulating the Conditional Posterior (ECP) algorithm, which uses emulation to increase the number of imputations, improving sampling efficiency for cut-posteriors. Mathews et al., (2025) designed a sequential Monte Carlo method to compute cut-posteriors.
@@
The existing variational inference approaches for cut-Bayes require access to the upstream raw data and model. These may not be available in many of the applications discussed above, where only the uncertainty quantified outputs are published (e.g., climate-model outputs, air pollution estimates, cosmological parameters, probabilistic word-embeddings). Sometimes the raw data may not be shared because of privacy concerns related to human-subjects research (e.g., uncertainty quantified estimates of cause-of-death classifiers published in Pramanik et al., 2025a, , are based on individual-level cause-of-death data that are not publicly available). Even if the upstream raw data and model are available, there is no need for the first variational approximation, as the cut-posterior for the upstream quantities is the same as their Bayes posterior from the upstream analysis. Also, many variational cut-Bayes methods rely on parametric assumptions for the variational family, such as the mean-field approximation that approximates marginal posteriors adequately but often fails to capture the complex conditional distributions. This is problematic, particularly when the true conditional cut-posterior deviates substantially from the chosen parametric form.
@@
In this manuscript, we propose a modular neural network-based variational inference approach for cut-Bayes that mitigates these limitations. Our contributions are twofold. First, we directly utilize samples of parameter distributions from the upstream analysis, and construct a single loss function that integrates all samples. When optimized, this estimates the variational conditional cut-posterior of the downstream parameters. This eliminates the need to access the upstream raw data and model, thereby enabling modularity of cut-Bayes analysis and protecting data privacy. Even if the upstream data and model are available, we simply use a single MCMC run to obtain upstream samples and use these in the loss for the downstream variational analysis. Thus, unlike the current approaches, we do not need a variational approximation for the upstream posterior thereby reducing additional approximation error. We use conditional normalizing flows, neural network-based conditional distribution families. to estimate the conditional cut-posterior of the downstream parameters. Our proposed method, Neural Variational Inference for Cut-Bayes (NeVI-Cut), thus leverages the expressive power of neural networks (Hornik,, 1991) and normalizing flows (Rezende and Mohamed,, 2015) enabling better approximation of complex, multi-modal, or skewed posteriors. Computationally, NeVI-Cut demonstrates significant improvement over methods like the multiple imputation-based nested MCMC for cutting feedback, which can also work only with upstream samples.
@@
A second contribution of this manuscript is developing a comprehensive theory for NeVI-Cut and variational cut-Bayes. Prior theoretical work on variational cut-Bayes is limited. Smith et al., (2025) prove consistency and asymptotic normality of the parameters of the Gaussian variational class for copula models, but their guarantees are asymptotic, appealing to Bernstein-von Mises type arguments that the true cut-posterior is approximately Gaussian asymptotically, thereby justifying the Gaussian variational family. Such results are reassuring yet do not address the quantity of real interest in applications, namely the quality of approximation to the exact cut-posterior conditional on the actually observed dataset, rather than a hypothetical large sample limit. On the other hand, Battaglia and Nicholls, (2024) provides a universal approximation result for conditional normalizing flows in terms of convergence in distribution. However, that is not sufficient to offer guarantees about the variational solutions using these flows, which minimize the Kullback-Leibler (KL) loss.
@@
Our theory is in the ‘fixed-data’ or ‘fixed-target’ regime, not asymptotic and not assuming any sample size growth. This is the more challenging regime of theory when studying the approximation quality of a variational solution to the target posterior. In an asymptotic paradigm, Bernstein-von Mises implies approximate Gaussian posteriors, so even simple Gaussian VI can be asymptotically adequate. That safety net is absent in the fixed target setting, where posteriors can have complicated shapes requiring sufficiently complex variational classes like normalizing flows. We keep the observed data fixed, and provide guarantees on the approximation quality of the variational estimate obtained from NeVI-Cut with respect to the cut-posterior conditional on the observed data. Our results explicitly quantify how the approximation rates depend on both the complexity of the target cut-posterior as well as the richness (number of parameters) of the neural architecture used to design the conditional flows. In the process, we establish, to our knowledge, novel results on universal and uniform approximation rates of conditional flows in the Kullback-Leibler divergence (KLD). Our results cover common conditional flow classes like Unconstrained Monotonic Neural Network flows (UMNN, Wehenkel and Louppe,, 2019) and neural Rational Quadratic Splines (RQS, Durkan et al.,, 2019). As the existing literature on flows and conditional flows (Huang et al.,, 2018; Papamakarios et al.,, 2021) has mostly focused on their representational capabilities, our results are of independent importance to study variational solutions using flows.
@@
The remainder of this paper is organized as follows: In Section 2, we illustrate the idea of cutting feedback. We then give background on variational inference and normalizing flows. In Section 3 we present our NeVI-Cut algorithm. Theoretical results are presented in Section 4. Section 5 details simulation experiments to illustrate the behavior of our model compared with the multiple imputation-based nested MCMC, the full Bayes approach, and parametric VI methods for cut-Bayes. Section 6 presents two real-world analyses using cutting feedback methods: a commonly used HPV data analysis, and the child and neonatal mortality data from the Countrywide Mortality Surveillance for Action (COMSA) Program in Mozambique. Section 7 concludes the paper with a discussion. Code to implement our proposed method is available on GitHub.
@@ 2.1 Cutting Feedback Approaches
We illustrate the concept of cutting feedback from downstream to upstream in Figure 1. Black arrows indicate parameters specifying a model, blue dotted arrows indicate the direction of information flow from a data source to parameters. Let $D_{1}$ denote the upstream data which informs the upstream quantity $\eta$. Here $\eta$ can be a set of parameters that enters the model for $D_{1}$ (in which case there would be a black arrow going from $\eta$ to $D_{1}$ and which is the standard cut model schematic used previously, e.g., in Plummer, (2015)). However, $\eta$ need not be a parameter in the model for $D_{1}$, it can be simply quantities that are informed by $D_{1}$ (e.g., derived quantities like predicted variables based on $D_{1}$ and its model). Our approach is agnostic to this. Hence, we do not specify a black arrow from $\eta$ to $D_{1}$ and simply show the eventual direction of information flow in the upstream analysis, that $\eta$ is informed by $D_{1}$. The downstream data is denoted by $D_{2}$ whose model involves both $\eta$ and some new parameters $\theta$ (hence the black arrows from $\eta$ and $\theta$ to $D_{2}$ in the left figure).
Figure 1: Paradigm of the cutting feedback from downstream to upstream.
@@
As we know in a Bayesian analysis, the final posterior of $\eta$ is informed by both $D_{1}$ and $D_{2}$ (hence the blue arrows from both $D_{1}$ and $D_{2}$ to $\eta$) and is thus different from its posterior $p(\eta\,|\,D_{1})$ after the upstream analysis. As discussed in the introduction, this is undesirable in many situations. For example, if the second module is misspecified (which is the case in many of the applications we highlighted), inference on $\eta$ could be negatively affected from this backward information flow from $D_{2}$ to $\eta$.
Cutting feedback from $D_{2}$ to $\eta$ (as indicated by the red dotted line) prevents this. In other words, instead of considering the Bayes posterior $p(\eta\,|\,D_{1},D_{2})$, we modify it to propagate the uncertainty in $\eta$ into the inference of $\theta$ from $D_{2}$ without feedback from $D_{2}$ to $\eta$. The marginal cut-posterior for $\eta$ is thus defined $p_{\text{cut}}(\eta\,|\,D_{1},D_{2})=p(\eta\,|\,D_{1})$, i.e., the posterior remains unchanged from the upstream analysis. We can then get the following joint cut-posterior:
$p_{\text{cut}}(\theta,\eta\,|\,D_{1},D_{2})$ $=\,p(\theta\,|\,\eta,D_{1},D_{2})\,p_{\text{cut}}(\eta\,|\,D_{1},D_{2})$ (2)
$=p(\theta\,|\,\eta,D_{2})\,p(\eta\,|\,D_{1}).$
Here, $p(\theta\,|\,\eta,D_{1},D_{2})=p(\theta\,|\,\eta,D_{2})$, as in (1), because the likelihood of $D_{1}$ does not depend of $\theta$ (hence no arrows between them), and as mentioned above, $p_{\text{cut}}(\eta\,|\,D_{1},D_{2})=p(\eta\,|\,D_{1})$ preserves the posterior of $\eta$ from upstream analysis and cuts the feedback from downstream. The marginal cut-posterior for $\theta$ is then given by
$p_{\text{cut}}(\theta\,|\,D_{1},D_{2})$ $=\int p(\theta\,|\,\eta,D_{2})\,p(\eta\,|\,D_{1})d\eta.$ (3)
Hence, to sample from the joint cut-posterior of $(\eta,\theta)$ or for the marginal cut-posterior of $\theta$, one can generate draws from $p(\theta\,|\,\eta,D_{2})$ for values of $\eta$ from $p(\eta\,|\,D_{1})$. This propagates the uncertainty of $\eta$ from the upstream analysis into the marginal cut-posterior for $\theta$. However, $\theta$ is typically multidimensional and direct sampling from $p(\theta\,|\,\eta,D_{2})$ is generally not feasible. Hence, for each draw of $\eta$ from $p(\eta\,|\,D_{1})$ (obtained from an MCMC run), a second nested MCMC is run to obtain samples of $\theta\,|\,\eta,D_{2}$. This nested MCMC scheme to sample from the cut-posterior is referred to as multiple imputation (MI; Plummer,, 2015) and is very computationally intensive as it requires MCMC runs for each sample for $\eta$ from the upstream analysis.

### v2

@@ Abstract
In many scientific applications, uncertainty of estimates from an earlier (upstream) analysis needs to be propagated to subsequent (downstream) analysis, without feedback (downstream data updating upstream posteriors). Cutting feedback methods, also termed cut-Bayes, achieve this by constructing a cut-posterior distribution. However, existing sampling-based approaches for cutting feedback, such as nested Markov chain Monte Carlo (MCMC), are computationally demanding, while variational inference (VI) methods require two separate variational approximations and access to the upstream data and model, which is often impractical in many applications due to privacy constraints or limited data accessibility. We propose NeVI-Cut, a provably accurate and modular neural network-based variational inference method for cutting feedback. We directly utilize samples from the upstream analysis without requiring access to the upstream data or model, and specify the conditional variational family for the downstream parameters using normalizing flows (neural network-based generative models). We provide fixed-data (not asymptotic) convergence rates of the NeVI-Cut solution in terms of richness of the neural architecture and the complexity of the exact target cut-posterior. In the process, we establish general results of independent importance on uniform Kullback-Leibler approximation rates of conditional distributions by common flow classes including Unconstrained Monotonic Neural Network and Neural Spline Flows. A triply stochastic algorithm implements the method efficiently. Simulation studies and two real-world analyses illustrate the speed and accuracy of NeVI-Cut.
Keywords: Bayesian analysis, Cutting Feedback, Variational Inference, Normalizing Flows, Neural Networks, Universal approximation rates.
@@ 1 Introduction
Across many scientific fields, uncertainty estimates of quantities from an upstream analysis are prestored to be reused in many downstream analyses. For example, downstream climate impacts studies (e.g., Butler et al.,, 2022; Park et al.,, 2023; Rao et al.,, 2024) commonly use ensemble outputs quantifying uncertainty from climate models such as CMIP6 (Thrasher et al.,, 2022). Similarly, the Global Burden of Disease project releases fine particulate matter (PM2.5) exposure and risk estimates as $1000$ posterior draws (Institute for Health Metrics and Evaluation,, 2024) that are used downstream in attributable-burden calculations (Murray et al.,, 2020). Importance of incorporating exposure uncertainty in downstream health outcomes analysis has been routinely emphasized (Peng and Bell,, 2010; Chang et al.,, 2011; Spiegelman,, 2016; Comess et al.,, 2024). Other examples of uncertainty quantified prestored outputs include posterior distributions of cosmological parameters, termed as Planck MCMC chains, posterior samples of confusion matrices of cause-of-death classifiers (Pramanik et al., 2025a, ) used for calibration of disease burden attribution, and probabilistic word-embeddings used for large language models (LLM) like word2Gauss and Bayesian skip-gram (Vilnis and McCallum,, 2014; Bražinskas et al.,, 2018) used in downstream text classification tasks.
@@
In many of the aforementioned applications, when propagating uncertainty to a downstream Bayesian analysis, it is desirable not to have feedback, which is the statistical phenomenon of the downstream data informing an updated distribution for the upstream parameters. This is because the upstream quantities are often more fundamental and causally precedent (e.g., air pollution concentrations or climate data) than downstream ones (e.g., individual-level health outcomes). More importantly, downstream models are often misspecified for the upstream quantities. For example, in downstream health association studies, upstream estimates of area-level air pollution outputs are often used only as proxies for the true unobserved individual-level exposures (Zeger et al.,, 2000). Similarly, confusion matrix estimates of classifiers trained on specific populations are often used in downstream calibration of different target populations (Pramanik et al., 2025a, ). Hence, it is undesirable to update these upstream parameter distributions based on the downstream data and model that are subject to model misspecifications, dataset shifts, and biases. Furthermore, many of these upstream quantities (e.g., climate outputs, air pollution estimates) are used in many different downstream analyses, and it is impractical to expect revised upstream posteriors to be published after every downstream study conducted by different groups.
@@
The idea of explicitly stopping downstream to upstream feedback is known as Cutting feedback (Liu et al.,, 2009; Plummer,, 2015; Jacob et al.,, 2017), which estimates the cut-posterior, a valid joint distribution over all parameters that preserves the distribution of the upstream parameters from the first analysis. There is now a large and rapidly growing literature on cutting feedback and related methods because of their broad applicability. A brief review is offered in Section S1 of the Supplement and Nott et al., (2023) offers a detailed overview. A major thread of developments has been computational. Plummer, (2015) showed that the ‘cut’ algorithm in the OpenBUGS package algorithm does not converge to the target cut-posterior. A principled alternative is the multiple imputation-based nested MCMC — running an MCMC for the upstream analysis to obtain a set of posterior samples, and then for each of these samples, running an MCMC for the downstream analysis, and then combining over all the MCMC posteriors. The repeated MCMC runs for the downstream analysis significantly increase computational cost. Plummer, (2015) proposed a tempered cut algorithm for approximating the cut-posterior. Jacob et al., (2020) developed an unbiased MCMC estimator using coupled chains and demonstrated the use for approximating the cut-posterior. Other notable work that has advanced the study of sampling based methods for cutting feedback include Liu and Goudie, (2022); Pompe and Jacob, (2021); Chakraborty et al., (2023) and others reviewed in Section S1.
@@
The existing variational inference approaches for cut-Bayes require access to the upstream raw data and model. These may not be available in many of the applications discussed above, where only the uncertainty quantified outputs are published (e.g., climate-model outputs, air pollution estimates, cosmological parameters, probabilistic word-embeddings). Sometimes the raw data may not be shared because of privacy concerns related to human-subjects research (e.g., uncertainty quantified estimates of cause-of-death classifiers published in Pramanik et al., 2025a, , are based on individual-level cause-of-death data that are not publicly available). Even if the upstream raw data and model are available, there is no need for the first variational approximation, as the cut-posterior for the upstream quantities is the same as their Bayes posterior from the upstream analysis. Also, many variational cut-Bayes methods rely on parametric assumptions for the variational family which is problematic when the true conditional cut-posterior deviates substantially from the chosen parametric form.
@@
In this manuscript, we propose a modular neural network-based variational inference approach for cut-Bayes that mitigates these limitations. Our contributions are twofold. First, we directly utilize samples of parameter distributions from the upstream analysis, and construct a single loss function that integrates all samples. When optimized, this estimates the variational conditional cut-posterior of the downstream parameters. This eliminates the need to access the upstream raw data and model, thereby enabling modularity of cut-Bayes analysis and protecting data privacy. Even if the upstream data and model are available, we simply use a single MCMC run to obtain upstream samples and use these in the loss for the downstream variational analysis. Thus, unlike the current approaches, we do not need a variational approximation for the upstream posterior thereby also reducing additional approximation error. We use conditional normalizing flows, neural network-based conditional distribution families to estimate the conditional cut-posterior of the downstream parameters. Our proposed method, Neural Variational Inference for Cut-Bayes (NeVI-Cut), thus leverages the expressive power of neural networks (Hornik,, 1991) and normalizing flows (Rezende and Mohamed,, 2015) enabling better approximation of complex, multi-modal, or skewed posteriors. Computationally, NeVI-Cut demonstrates significant improvement over sampling-based methods like the nested MCMC for cutting feedback, while being substantially more accurate than parametric variational cutting feedback methods.
@@
A second contribution of this manuscript is developing a comprehensive theory for NeVI-Cut and variational cut-Bayes. Prior theoretical work on variational cut-Bayes is limited. Smith et al., (2025) proves consistency and asymptotic normality of the parameters of the Gaussian variational class for copula models, but their guarantees are asymptotic, appealing to Bernstein-von Mises type arguments that the true cut-posterior is approximately Gaussian asymptotically, thereby justifying the Gaussian variational family. Such results are reassuring yet do not address the quantity of real interest in applications, namely the quality of approximation to the exact cut-posterior conditional on the actually observed dataset, rather than a hypothetical large sample limit. On the other hand, Battaglia and Nicholls, (2024) provides a universal approximation result for conditional normalizing flows in terms of convergence in distribution. However, that is not sufficient to offer guarantees about the variational solutions using these flows, which minimize the Kullback-Leibler divergence (KLD).
@@
Our theory is in the ‘fixed-data’ or ‘fixed-target’ regime, not asymptotic and not assuming any sample size growth. This is the more challenging regime of theory when studying the approximation quality of a variational solution to the target posterior. In an asymptotic paradigm, Bernstein-von Mises implies approximate Gaussian posteriors, so even simple Gaussian VI can be asymptotically adequate. That safety net is absent in the fixed target setting, where posteriors can have complicated shapes requiring sufficiently complex variational classes like normalizing flows. We keep the observed data fixed, and provide guarantees on the approximation quality of the variational estimate obtained from NeVI-Cut with respect to the cut-posterior conditional on the observed data. Our results explicitly quantify how the approximation rates depend on both the complexity of the target cut-posterior as well as the richness (number of parameters) of the neural architecture used to design the conditional flows. In the process, we establish, to our knowledge, novel results on universal and uniform approximation KLD rates of conditional flows. Our results cover common conditional flow classes like Unconstrained Monotonic Neural Network flows (UMNN; Wehenkel and Louppe,, 2019) and Rational Quadratic Neural Spline Flows (RQ-NSF; Durkan et al.,, 2019). As the existing literature on flows and conditional flows (Huang et al.,, 2018; Papamakarios et al.,, 2021) has mostly focused on their representational capabilities, our results are of independent importance, as they facilitate study of variational solutions using flows.
@@
The remainder of this paper is organized as follows: In Section 2, we illustrate the idea of cutting feedback. We then give background on variational inference and normalizing flows. In Section 3 we present our NeVI-Cut algorithm. Theoretical results are presented in Section 4. Section 5 details simulation experiments to benchmark speed and accuracy of our method. Section 6 presents two real-world analyses using cutting feedback methods: a commonly used HPV data analysis, and the child and neonatal mortality data from the Countrywide Mortality Surveillance for Action (COMSA) Program in Mozambique. Section 7 concludes the paper with a discussion, to illustrate benefits of NeVI-Cut. Code to implement our proposed method is available on NeVI-Cut Python package on GitHub.
@@ 2.1 Cutting Feedback Approaches
Let $D_{1}$ denote the upstream data which informs the upstream quantity (parameters or predictive variables) $\eta$. The downstream data is denoted by $D_{2}$ whose model involves both $\eta$ and some new parameters $\theta$. If both datasets were analyzed jointly, the Bayesian posterior distribution can be written as:
@@
In a Bayesian analysis, the final posterior of $\eta$ is informed by both $D_{1}$ and $D_{2}$ and is thus different from its posterior $p(\eta\,|\,D_{1})$ after the upstream analysis. As discussed in the Introduction, this is undesirable in many situations. Cutting feedback prevents this by targeting the cut-posterior
$p_{\text{cut}}(\theta,\eta\,|\,D_{1},D_{2})=\,p(\theta\,|\,\eta,D_{1},D_{2})\,p_{\text{cut}}(\eta\,|\,D_{1},D_{2})=p(\theta\,|\,\eta,D_{2})\,p(\eta\,|\,D_{1}).$ (2)
Here, $p(\theta\,|\,\eta,D_{1},D_{2})=p(\theta\,|\,\eta,D_{2})$, as in (S1), because the likelihood of $D_{1}$ does not depend of $\theta$ (hence no arrows between them), and as mentioned above, $p_{\text{cut}}(\eta\,|\,D_{1},D_{2})=p(\eta\,|\,D_{1})$ preserves the posterior of $\eta$ from upstream analysis. A more detailed introduction of cutting feedback is offered in Section S1.